OpenStudy (anonymous):

can (25h-8)/((3h)(h-2)) be simplified?

OpenStudy (anonymous):

doesn't look like it to me...

OpenStudy (anonymous):

\[\frac{ 25h-8 }{ 3h ^{2}-6h }\]

OpenStudy (anonymous):

then i factored out a 3h..

OpenStudy (anonymous):

yeah I don't see much more you can do with it.

OpenStudy (anonymous):

me either

OpenStudy (anonymous):

so i got (25h-8)/3h(h-2)

OpenStudy (anonymous):

i think that's as good as it gets

OpenStudy (anonymous):

but that' s still not the right answer.... my program isn't taking it

OpenStudy (anonymous):

\[\frac{ 25 }{3h-6 } - \frac{ 8 }{ 3h ^{2} -6h}\]

OpenStudy (anonymous):

this is an expression though so you can't multiply by 3h

OpenStudy (anonymous):

How did you get that? and You can multiply by 3h/3h which is 1

OpenStudy (anonymous):

but then you'd be back at the same expression

OpenStudy (anonymous):

actually I see what your saying you then want to take out 1/3h right

OpenStudy (anonymous):

the original problem is (21h+4h-8)/((3h)(h-2))

OpenStudy (anonymous):

why didnt you post that from the beginning :/

OpenStudy (anonymous):

Ya, I deleted my post because I think it was a dead end

OpenStudy (anonymous):

so that gives you (25h-8)/((3h)(h-2))

OpenStudy (anonymous):

oh I think I see an answer

OpenStudy (anonymous):

There must be a reason 21h and 4h is split up.

OpenStudy (anonymous):

you need to get h-2 through factorization

OpenStudy (anonymous):

\[\frac{ 21h +4(h-2) }{3h(h-2) }\]

OpenStudy (anonymous):

\[\frac{ 21h + (h-2)(4) }{ 3h(h-2) }\]

OpenStudy (anonymous):

haha

OpenStudy (anonymous):

split up into 2 different fractions next

OpenStudy (anonymous):

that one got me

OpenStudy (anonymous):

now try to solve it mberdeja

OpenStudy (anonymous):

so it's (21+4)/(3h)

OpenStudy (anonymous):

(21h+4)/(3h)

OpenStudy (anonymous):

no you have to break up the fraction.

OpenStudy (anonymous):

\[\frac{ 21 }{ 3h-6 } + \frac{ 4 }{ 3h }\]

OpenStudy (anonymous):

no cause the h-2 cancels at the bottom and at the top

OpenStudy (anonymous):

your right but if you look the top is conjoined by an addition operator

OpenStudy (anonymous):

you cant just cancel terms unless the term is multiplied by the entire thing which it isn't